# Compositional Monte Carlo Tree Diffusion for Extendable Planning

### v1

@@ 5.4 High-dimensional visual navigation: POMDP planning challenges
We evaluate C-MCTD on Visual Maze tasks from prior work [32], which require navigation from starting observations to target goals within $64\times 64$ RGB observation spaces—forming challenging Partially Observable Markov Decision Processes (POMDPs). Following established protocols [32], we employ Variational Autoencoders [18] for latent space planning, inverse dynamics models for action generation, and positional estimators for POMDP guidance (details in Appendix A.5.3).
@@ 7 Conclusion
We would like to extend our gratitude to Doojin Baek for his insightful discussions and assistance throughout this project. This research was supported by Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT (MSIT) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509279, Global AI Frontier Lab).
@@ A.2 Environment details
To evaluate our C-MCTD, we adopt the benchmarks from Yoon et al. [32]. While the robotic manipulation and visual maze tasks replicate the environmental settings of the prior work, we introduce a more demanding evaluation for the long-horizon maze environments. Specifically, we employ the stitch datasets from Park et al. [28] rather than the navigate datasets. This choice creates a challenging scenario that requires generating plans significantly longer than the trajectories observed during training. For instance, training trajectories are limited to 200 steps, whereas successfully reaching goals in the test environments requires plans of up to 1000 steps. This setting allows us to rigorously assess the extendable planning capabilities of our approach. An overview of each task is provided in Figure 3.
@@ A.3.1 Stitching methods
[6, 21] An alternative approach, which we term DataStitch, extends the planning horizon by first elongating the trajectories within the training dataset itself. Because trajectories in the original dataset are not necessarily connectable, this method involves training auxiliary diffusion planners and inverse dynamics models to generate synthetic continuations, effectively stitching existing data points together. For instance, Chen et al. [6] iteratively applied this process to extend trajectories up to sevenfold. Following this methodology, we create an extended version of the stitch dataset by lengthening trajectories to five times their original length and subsequently train our models on this augmented data.
@@ A.3.2 Diffusion planners
[15] As a primary baseline, we include Diffuser, which pioneered training diffusion models on sequences of concatenated state-action pairs. We evaluate Diffuser with both the Replan and DataStitch extension methods by training it on the corresponding augmented datasets.
@@
[2] This Diffuser variant tokenizes data and applies varied noise levels to enable causal denoising that prioritizes near-future states. We employ the Transformer-based implementation proposed by Chen et al. [2]. As the method is inherently designed for replanning, we evaluate it with the DataStitch extension but do not create a separate “Replan” variant.
@@
[17] Kim et al. [17] introduced a diffusion planner that achieves extendability via a learned value function, which guides the model to concatenate short trajectories into a longer, task-relevant plan. While effective, we empirically found that learning a reliable value function is challenging with suboptimal data, whereas inference-time scaling offers a more robust alternative. Because SSD is natively designed for extendable planning, we do not evaluate it using our Replan or DataStitch extensions.
@@
[24] Luo et al. [24] proposed an alternative approach to extendable planning where a diffusion model is trained to generate sub-trajectories conditioned on nearby context. This framework allows the model to compose multiple sub-trajectories, generating plans that are longer than any seen during training. Similar to SSD, CompDiffuser’s native extendability makes evaluation with our proposed extensions unnecessary.
@@
[32] Yoon et al. [32] introduced a framework that implements inference-time scaling by branching the denoising process into a Monte Carlo Tree Search, rather than merely increasing sequential denoising steps. Guided by a reward function, this approach explores alternative paths to find optimal solutions, demonstrating superior performance across diverse planning tasks. We evaluate MCTD with both the Replan and DataStitch extension methods.
@@ A.4 Model hyperparameters
Our hyperparameter configuration is largely adopted from Yoon et al. [32]. For full reproducibility, we provide the detailed configurations for the baseline models in Tables 58. Hyperparameters introduced by our proposed C-MCTD are detailed in Section A.5 alongside their corresponding experimental setups. For the SSD baseline, we utilized the default configurations from its official public implementation [^1] https://github.com/rlatjddbs/SSD.
Table 5: Hyperparameters for the Diffuser baseline. These settings are adopted from the original MCTD paper [32] to ensure a fair comparison. Task-specific parameters, such as the planning horizon, are detailed in their respective sections.
@@
Table 6: Hyperparameters for the Diffusion Forcing baseline. Settings are based on the implementation in Yoon et al. [32] and adapted for our tasks. Parameters that vary across environments, such as the guidance scale, are specified in Section A.5.
@@
Table 7: Hyperparameters for the Monte Carlo Tree Diffusion (MCTD) baseline. These settings are based on the original implementation [32]. Task-dependent parameters, such as the planning horizon and guidance set, are specified in Section A.5.
@@ A.5.1 Long-horizon maze environments
To evaluate extendable planning, all models were trained on the challenging stitch datasets. An exception was made for the value-learning policy [31], which failed to converge on this data; as this baseline is not our focus, we trained it on the original navigate dataset. For all maze environments, we set the training trajectory length and the single-pass planning horizon to 100, sampling sub-sequences from the full 200-step trajectories to enhance data diversity. Following prior work [32], we remove random noise from the start and goal positions to isolate the planner’s performance from environmental stochasticity. Furthermore, we employ a heuristic controller for PointMaze environments [15] and a value-based policy for AntMaze environments [31, 4], for which we set the subgoal interval to $10$.
@@
  • Diffusion Forcing: Following Chen et al. [2], we set the guidance scale to 3 for medium maps and 2 for large and giant maps.
  • MCTD: The guidance sets are adopted from Yoon et al. [32]: $\{0,0.1,0.5,1,2\}$ for PointMaze (Medium & Large), $\{0.5,1,2,3,4\}$ for PointMaze (Giant), and $\{0,1,2,3,4,5\}$ for all AntMaze tasks.
@@
Both MCTD and C-MCTD require a reward function to evaluate simulated trajectories. We adopt the function from Yoon et al. [32], which incorporates two components. First, it penalizes physically implausible plans by assigning a reward of 0 to any trajectory with unrealistic position changes between consecutive states. Second, for valid trajectories that first reach the goal at timestep $t$, the reward is calculated as r=(Ht)/Hr=(H-t)/H, where $H$ is the maximum horizon length. This encourages finding shorter, more efficient paths.
@@
  • Node expansion planner: For node expansion in the maze environments, C-MCTD utilizes a best-of-$N$ search-based planner [34] built upon Diffusion Forcing [2]. In this approach, $N$ candidate trajectories are generated, and the one yielding the highest reward (as defined above) is selected. We set $N=50$ for all maze experiments.
  • DC and PC variants: Our Distributed Composer (DC) and Preplan Composer (PC) variants identify representative waypoints via k-means clustering [22, 26] on the training data, with 10, 30, and 70 cluster centers for medium, large, and giant maps, respectively. For DC, the search is capped at 100 iterations. For PC, we limit the search to 20 iterations and 2 plan concatenations (a tree depth of 2) for medium and large maps; these are reduced to 10 iterations and 1 concatenation for the giant map. These settings render our methods substantially more efficient than the Online Composer baseline, which requires up to 500 search iterations and 10 concatenations.
@@ A.5.2 Robot arm manipulation environment
We conduct these experiments using the play datasets. The planning horizon is set to 200 for single-cube tasks, corresponding to the maximum episode length, and extended to 500 for more complex multi-cube tasks. Similar to the AntMaze experiments, we employ a value-based policy [31] as a low-level controller, with a subgoal interval of 10.
@@
The goal-reaching guidance function is identical to that used in the maze environments (Section A.5.1). For the Diffusion Forcing baseline, the guidance scale is set to 2. For tree-search planners (MCTD and C-MCTD), we adopt the object-wise guidance strategy from Yoon et al. [32]. This approach guides the planner to manipulate a single cube at a time, preventing physically infeasible plans that attempt to move multiple objects concurrently. The guidance set for this strategy is $\{1,2,4\}$, which our C-MCTD variants inherit for two child nodes during expansion.
@@
We adopt the reward function from Yoon et al. [32], which is designed to filter out physically unrealistic trajectories. A plan receives a reward of 0 if it violates any of the following conditions: (1) moving multiple objects simultaneously; (2) leaving an object in an unstable, mid-air position; (3) causing inter-object collisions; or (4) attempting to grasp an object obstructed by another. Otherwise, a positive reward is assigned based on task success.
@@
Following Yoon et al. [32], we incorporate a set of pre-defined macro-actions (scripted primitives) to improve the efficiency and robustness of object-wise planning. These routines handle common operations, such as automatically releasing an object at its target or pre-positioning the arm before manipulating the next object. This scripting approach prevents common planning failures, like generating trajectories beyond the arm’s kinematic reach or attempting to grasp with an occupied gripper.
@@ A.5.3 Visual maze environment
We follow the experimental setup from Yoon et al. [32] for the visual maze tasks. Using the data generation scripts from Park et al. [28], we create datasets of 1000-step trajectories for both medium and large maps. To ensure diverse training samples, our diffusion models are trained on sub-trajectories with a planning horizon of 500, sampled from these full-length trajectories.
@@
To handle the high-dimensional visual inputs and partial observability, we adopt the modeling pipeline from Yoon et al. [32], which consists of three pre-trained components that operate on $64\times 64\times 3$ RGB observations:
  • Variational Autoencoder (VAE): A VAE [18] is first pre-trained to encode each image observation into a compact 8-dimensional latent representation, $z_{t}$. This latent space serves as the basis for all subsequent planning and control models.
@@
Guidance is performed in the state space using the positional signal from the pre-trained position estimator. This allows us to apply the same L2 distance-based guidance function as described in Appendix A.5.1. Following Yoon et al. [32], we use a guidance set of $\{0,0.1,0.5,1,2\}$ for all baselines. Our C-MCTD variants inherit this configuration, assigning this guidance set to each child node during an expansion.
@@ B.3 Ablation Study on Stitching Threshold
We investigate the sensitivity of C-MCTD to the stitching threshold hyperparameter, $\epsilon$. In our main experiments, we set $\epsilon=1.0$, adopting the value from the goal-achievement threshold in the original MCTD framework [32]. To analyze the robustness of our method to this choice, we conducted an extensive ablation study with our Distributed Composer.

### v2

@@ 5.4 High-dimensional visual navigation: POMDP planning challenges
We evaluate C-MCTD on Visual Maze tasks from prior work [mctd], which require navigation from starting observations to target goals within $64\times 64$ RGB observation spaces—forming challenging Partially Observable Markov Decision Processes (POMDPs). Following established protocols [mctd], we employ Variational Autoencoders [vae] for latent space planning, inverse dynamics models for action generation, and positional estimators for POMDP guidance (details in Appendix A.5.3).
@@ 7 Conclusion
We would like to extend our gratitude to Doojin Baek for his insightful discussions and assistance throughout this project. This research was supported by Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT (MSIT), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00509279, Global AI Frontier Lab), and GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT (MSIT).
@@ A.2 Environment details
To evaluate our C-MCTD, we adopt the benchmarks from mctd. While the robotic manipulation and visual maze tasks replicate the environmental settings of the prior work, we introduce a more demanding evaluation for the long-horizon maze environments. Specifically, we employ the stitch datasets from ogbench rather than the navigate datasets. This choice creates a challenging scenario that requires generating plans significantly longer than the trajectories observed during training. For instance, training trajectories are limited to 200 steps, whereas successfully reaching goals in the test environments requires plans of up to 1000 steps. This setting allows us to rigorously assess the extendable planning capabilities of our approach. An overview of each task is provided in Figure 3.
@@ A.3.1 Stitching methods
[hmd, diffstitch] An alternative approach, which we term DataStitch, extends the planning horizon by first elongating the trajectories within the training dataset itself. Because trajectories in the original dataset are not necessarily connectable, this method involves training auxiliary diffusion planners and inverse dynamics models to generate synthetic continuations, effectively stitching existing data points together. For instance, hmd iteratively applied this process to extend trajectories up to sevenfold. Following this methodology, we create an extended version of the stitch dataset by lengthening trajectories to five times their original length and subsequently train our models on this augmented data.
@@ A.3.2 Diffusion planners
[diffuser] As a primary baseline, we include Diffuser, which pioneered training diffusion models on sequences of concatenated state-action pairs. We evaluate Diffuser with both the Replan and DataStitch extension methods by training it on the corresponding augmented datasets.
@@
[diffusion_forcing] This Diffuser variant tokenizes data and applies varied noise levels to enable causal denoising that prioritizes near-future states. We employ the Transformer-based implementation proposed by diffusion_forcing. As the method is inherently designed for replanning, we evaluate it with the DataStitch extension but do not create a separate “Replan” variant.
@@
[ssd] ssd introduced a diffusion planner that achieves extendability via a learned value function, which guides the model to concatenate short trajectories into a longer, task-relevant plan. While effective, we empirically found that learning a reliable value function is challenging with suboptimal data, whereas inference-time scaling offers a more robust alternative. Because SSD is natively designed for extendable planning, we do not evaluate it using our Replan or DataStitch extensions.
@@
[compdiffuser] compdiffuser proposed an alternative approach to extendable planning where a diffusion model is trained to generate sub-trajectories conditioned on nearby context. This framework allows the model to compose multiple sub-trajectories, generating plans that are longer than any seen during training. Similar to SSD, CompDiffuser’s native extendability makes evaluation with our proposed extensions unnecessary.
@@
[mctd] mctd introduced a framework that implements inference-time scaling by branching the denoising process into a Monte Carlo Tree Search, rather than merely increasing sequential denoising steps. Guided by a reward function, this approach explores alternative paths to find optimal solutions, demonstrating superior performance across diverse planning tasks. We evaluate MCTD with both the Replan and DataStitch extension methods.
@@ A.4 Model hyperparameters
Our hyperparameter configuration is largely adopted from mctd. For full reproducibility, we provide the detailed configurations for the baseline models in Tables 58. Hyperparameters introduced by our proposed C-MCTD are detailed in Section A.5 alongside their corresponding experimental setups. For the SSD baseline, we utilized the default configurations from its official public implementation [^1] https://github.com/rlatjddbs/SSD.
Table 5: Hyperparameters for the Diffuser baseline. These settings are adopted from the original MCTD paper [mctd] to ensure a fair comparison. Task-specific parameters, such as the planning horizon, are detailed in their respective sections.
@@
Table 6: Hyperparameters for the Diffusion Forcing baseline. Settings are based on the implementation in mctd and adapted for our tasks. Parameters that vary across environments, such as the guidance scale, are specified in Section A.5.
@@
Table 7: Hyperparameters for the Monte Carlo Tree Diffusion (MCTD) baseline. These settings are based on the original implementation [mctd]. Task-dependent parameters, such as the planning horizon and guidance set, are specified in Section A.5.
@@ A.5.1 Long-horizon maze environments
To evaluate extendable planning, all models were trained on the challenging stitch datasets. An exception was made for the value-learning policy [dql], which failed to converge on this data; as this baseline is not our focus, we trained it on the original navigate dataset. For all maze environments, we set the training trajectory length and the single-pass planning horizon to 100, sampling sub-sequences from the full 200-step trajectories to enhance data diversity. Following prior work [mctd], we remove random noise from the start and goal positions to isolate the planner’s performance from environmental stochasticity. Furthermore, we employ a heuristic controller for PointMaze environments [diffuser] and a value-based policy for AntMaze environments [dql, plandq], for which we set the subgoal interval to $10$.
@@
  • Diffusion Forcing: Following diffusion_forcing, we set the guidance scale to 3 for medium maps and 2 for large and giant maps.
  • MCTD: The guidance sets are adopted from mctd: $\{0,0.1,0.5,1,2\}$ for PointMaze (Medium & Large), $\{0.5,1,2,3,4\}$ for PointMaze (Giant), and $\{0,1,2,3,4,5\}$ for all AntMaze tasks.
@@
Both MCTD and C-MCTD require a reward function to evaluate simulated trajectories. We adopt the function from mctd, which incorporates two components. First, it penalizes physically implausible plans by assigning a reward of 0 to any trajectory with unrealistic position changes between consecutive states. Second, for valid trajectories that first reach the goal at timestep $t$, the reward is calculated as r=(Ht)/Hr=(H-t)/H, where $H$ is the maximum horizon length. This encourages finding shorter, more efficient paths.
@@
  • Node expansion planner: For node expansion in the maze environments, C-MCTD utilizes a best-of-$N$ search-based planner [d-mpc] built upon Diffusion Forcing [diffusion_forcing]. In this approach, $N$ candidate trajectories are generated, and the one yielding the highest reward (as defined above) is selected. We set $N=50$ for all maze experiments.
  • DC and PC variants: Our Distributed Composer (DC) and Preplan Composer (PC) variants identify representative waypoints via k-means clustering [kmeans1, kmeans2] on the training data, with 10, 30, and 70 cluster centers for medium, large, and giant maps, respectively. For DC, the search is capped at 100 iterations. For PC, we limit the search to 20 iterations and 2 plan concatenations (a tree depth of 2) for medium and large maps; these are reduced to 10 iterations and 1 concatenation for the giant map. These settings render our methods substantially more efficient than the Online Composer baseline, which requires up to 500 search iterations and 10 concatenations.
@@ A.5.2 Robot arm manipulation environment
We conduct these experiments using the play datasets. The planning horizon is set to 200 for single-cube tasks, corresponding to the maximum episode length, and extended to 500 for more complex multi-cube tasks. Similar to the AntMaze experiments, we employ a value-based policy [dql] as a low-level controller, with a subgoal interval of 10.
@@
The goal-reaching guidance function is identical to that used in the maze environments (Section A.5.1). For the Diffusion Forcing baseline, the guidance scale is set to 2. For tree-search planners (MCTD and C-MCTD), we adopt the object-wise guidance strategy from mctd. This approach guides the planner to manipulate a single cube at a time, preventing physically infeasible plans that attempt to move multiple objects concurrently. The guidance set for this strategy is $\{1,2,4\}$, which our C-MCTD variants inherit for two child nodes during expansion.
@@
We adopt the reward function from mctd, which is designed to filter out physically unrealistic trajectories. A plan receives a reward of 0 if it violates any of the following conditions: (1) moving multiple objects simultaneously; (2) leaving an object in an unstable, mid-air position; (3) causing inter-object collisions; or (4) attempting to grasp an object obstructed by another. Otherwise, a positive reward is assigned based on task success.
@@
Following mctd, we incorporate a set of pre-defined macro-actions (scripted primitives) to improve the efficiency and robustness of object-wise planning. These routines handle common operations, such as automatically releasing an object at its target or pre-positioning the arm before manipulating the next object. This scripting approach prevents common planning failures, like generating trajectories beyond the arm’s kinematic reach or attempting to grasp with an occupied gripper.
@@ A.5.3 Visual maze environment
We follow the experimental setup from mctd for the visual maze tasks. Using the data generation scripts from ogbench, we create datasets of 1000-step trajectories for both medium and large maps. To ensure diverse training samples, our diffusion models are trained on sub-trajectories with a planning horizon of 500, sampled from these full-length trajectories.
@@
To handle the high-dimensional visual inputs and partial observability, we adopt the modeling pipeline from mctd, which consists of three pre-trained components that operate on $64\times 64\times 3$ RGB observations:
  • Variational Autoencoder (VAE): A VAE [vae] is first pre-trained to encode each image observation into a compact 8-dimensional latent representation, $z_{t}$. This latent space serves as the basis for all subsequent planning and control models.
@@
Guidance is performed in the state space using the positional signal from the pre-trained position estimator. This allows us to apply the same L2 distance-based guidance function as described in Appendix A.5.1. Following mctd, we use a guidance set of $\{0,0.1,0.5,1,2\}$ for all baselines. Our C-MCTD variants inherit this configuration, assigning this guidance set to each child node during an expansion.
@@ B.3 Ablation Study on Stitching Threshold
We investigate the sensitivity of C-MCTD to the stitching threshold hyperparameter, $\epsilon$. In our main experiments, we set $\epsilon=1.0$, adopting the value from the goal-achievement threshold in the original MCTD framework [mctd]. To analyze the robustness of our method to this choice, we conducted an extensive ablation study with our Distributed Composer.